OpenStudy (anonymous):

How do I find the trace of the curve gama (t) = (e^(-t) cos t, e^(-t) sin t)?

OpenStudy (anonymous):



OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

\[\gamma \left( t \right) = (e ^{-t} \cos {t} , e^{-t} \sin {t})\]